15.3 Consider the following $1,000 par value zero-coupon bonds: Bond Years to Maturity YTM(%) A 1...

15.3

Consider the following $1,000 par value zero-coupon bonds: Bond Years to Maturity YTM(%) A 1 5.7 % B 2 6.7 C 3 7.2 D 4 7.7 According to the expectations hypothesis, what is the market’s expectation of the yield curve one year from now? Specifically, what are the expected values of next year’s yields on bonds with maturities of (a) one year? (b) two years? (c) three years? (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

1.
=(1+6.7%)^2/(1+5.7%)^1-1
=7.7095%

2.
=(1+7.2%)^3/(1+6.7%)^2-1
=8.2070%

3.
=(1+7.7%)^4/(1+7.2%)^3-1
=9.2140%
Answer #2

(a). ((1+6.7%)^2/(1+5.7%)^1)^(1/1)-1


(b). ((1+7.2%)^3/(1+5.7%)^1)^(1/2)-1


(c). ((1+7.7%)^4/(1+5.7%)^1)^(1/3)-1

answered by: Egg